OpenStudy (anonymous):

There are 22 students in Mario's class. Mario's teacher is planning to pair the students up for a project. How many different combinations of pairs are possible?

OpenStudy (anonymous):

\[\frac{ 22! }{ (22-2) !}\]

OpenStudy (luigi0210):

\[\frac{ n! }{ (n-r)!r! }\]

OpenStudy (anonymous):

11 44 231 462

OpenStudy (anonymous):

Thanks! Mind helping with some more?

OpenStudy (luigi0210):

This is combinations not permutations peru..

OpenStudy (anonymous):

Okay so what is the answer then?

OpenStudy (whpalmer4):

\[\frac{22!}{(22-2)! 2!} = \frac{22*21*\cancel{20*19*18...3*2*1}}{\cancel{(20*19*18*...3*2*1)}(2*1)} = \frac{22*21}{2} = 231\]